OpenStudy (anonymous):

1. Excel Simulation #1: Say that every single class in Full Sails Masters Degree in Game Design has a 5% failure rate per month. Say also that when a student fails once, he or she drops from the program. There are twelve months. Given a cohort of 100 that starts the program, about how many are left after twelve months? Remember that someone can only pass or fail.

OpenStudy (fibonaccichick666):

well if you start with 100 students, and you lose 5% in the first month that leaves you with?

OpenStudy (anonymous):

95, since 5% of 100 is 5 correct?

OpenStudy (fibonaccichick666):

yep now to start month 2 you have 95 people, you lose another 5% how many are left?

OpenStudy (fibonaccichick666):

remember you cannot have a partial person, so you need to round up

OpenStudy (anonymous):

90 people left?

OpenStudy (fibonaccichick666):

correct because \[0.05*95=4.75\]\[95-5=90\] now can you continue this process?

OpenStudy (anonymous):

Indeed I can thank you for the help. Now to trying and work it all together on Excel

OpenStudy (anonymous):

And thank you*

OpenStudy (fibonaccichick666):

oh on excel its just a formula and np if you need excel I can do that to

OpenStudy (anonymous):

That would be awesome, first time using formulas on excel so am a tad lost

OpenStudy (fibonaccichick666):

ok np so I'm going to draw up a doc real quick so I can run you through it

OpenStudy (anonymous):

Thank you very much

OpenStudy (fibonaccichick666):

Alright so sorry about the wait, I can't attah the file because my comp keeps crashing. But take column A and in the second row label that month

OpenStudy (anonymous):

got it

OpenStudy (fibonaccichick666):

now in B1 label that total students

OpenStudy (anonymous):

alright

OpenStudy (fibonaccichick666):

now in A3 put a 0 and in A4 put a 1 then use the autofill handle down to 12

OpenStudy (fibonaccichick666):

so you have 12 total months

OpenStudy (anonymous):

done

OpenStudy (fibonaccichick666):

now we want to write that process we used to solve the equation in math terms

OpenStudy (fibonaccichick666):

so if x=total students y=students after first drop out what is our eq?

OpenStudy (anonymous):

y=X-0.05?

OpenStudy (fibonaccichick666):

well that is almost it but remember .05 times what?

OpenStudy (anonymous):

100. So 0.05*100 = 5

OpenStudy (fibonaccichick666):

right but 100 = what in our variables?

OpenStudy (anonymous):

total starting students

OpenStudy (fibonaccichick666):

or what letter?

OpenStudy (anonymous):

X

OpenStudy (fibonaccichick666):

right so can you rewrite the formula?

OpenStudy (anonymous):

y=100-5?

OpenStudy (fibonaccichick666):

no use your previous eq that you had written

OpenStudy (anonymous):

y=100-0.05

OpenStudy (fibonaccichick666):

you had y=x-.05

OpenStudy (anonymous):

y=100-0.05

OpenStudy (fibonaccichick666):

no keep it in terms of x

OpenStudy (fibonaccichick666):

if 5=.05(100) then that is the same as saying 5=.05(X)

OpenStudy (fibonaccichick666):

so y =?

OpenStudy (anonymous):

haha got it now

OpenStudy (fibonaccichick666):

ok so what is the formula?

OpenStudy (anonymous):

haha ive been at this for so long today that am brain dead I swear. y=95

OpenStudy (fibonaccichick666):

in terms of x please

OpenStudy (anonymous):

95=100-.05

OpenStudy (fibonaccichick666):

in terms of x!

OpenStudy (fibonaccichick666):

variables only no numbers

OpenStudy (fibonaccichick666):

btw what is 100-.05 and does that =95?

OpenStudy (anonymous):

I honestly don't know how to write the formula out. Maybe if i see it i can come to sense of it then resume to punch myself in the face cause it was so easy

OpenStudy (fibonaccichick666):

you already wrote it! lol just remember that .05x not .05

OpenStudy (anonymous):

y=.05x

OpenStudy (fibonaccichick666):

nooooo! ugh ok let's go through this again. what is y?

OpenStudy (fibonaccichick666):

sorry i'm just frustrated I shouldn't have "ughhh-ed" you

OpenStudy (fibonaccichick666):

I mean what does y mean

OpenStudy (fibonaccichick666):

our definition of y was?

OpenStudy (anonymous):

AHHH sigh... Alright so Y= students after first drop out

OpenStudy (fibonaccichick666):

good now what is x?

OpenStudy (anonymous):

X=total of students

OpenStudy (fibonaccichick666):

now how did we get y when we were doing numbers? (in words)

OpenStudy (anonymous):

Multiplied 0.05 by 100 (total number of students atm) to get 5 then subtract 5 by 100

OpenStudy (anonymous):

to get 95

OpenStudy (fibonaccichick666):

now write the first part with x=100

OpenStudy (anonymous):

.05(100)

OpenStudy (fibonaccichick666):

in one eq write the entire sentence

OpenStudy (fibonaccichick666):

in terms of x

OpenStudy (fibonaccichick666):

do NOT write 100

OpenStudy (anonymous):

.05(x)

OpenStudy (fibonaccichick666):

that only gives you how many people dropout. I want the total left

OpenStudy (anonymous):

.05(x)-X=y

OpenStudy (fibonaccichick666):

that would be a negative number

OpenStudy (anonymous):

y=X-.05(x)

OpenStudy (fibonaccichick666):

YES!

OpenStudy (fibonaccichick666):

thats it!

OpenStudy (anonymous):

Punchinh my self in the face atm

OpenStudy (anonymous):

lmao

OpenStudy (fibonaccichick666):

lol ok now, the x you have will be the cell above

OpenStudy (fibonaccichick666):

at month 0 you have 100 insert the formula starting with month 1

OpenStudy (fibonaccichick666):

then copy the formula down

OpenStudy (anonymous):

how will i got about inserting it since Excel on read =PRODUCT for example

OpenStudy (anonymous):

only reads*

OpenStudy (fibonaccichick666):

no... you just type what i told you only instead of x click on the 100 cell should be like B2

OpenStudy (fibonaccichick666):

there is no formula other than what you already have you don't need an excel fn

OpenStudy (anonymous):

go it

OpenStudy (anonymous):

got it*

OpenStudy (fibonaccichick666):

ok!!! yay congrats

OpenStudy (anonymous):

haha thank you very much

OpenStudy (fibonaccichick666):

just remember to format the cells to 00 decimal places so you only have whole numbers

OpenStudy (anonymous):

Thank you very much

OpenStudy (fibonaccichick666):

Np, if you think I did a good job-and you have time- could you just write a testimonial?

OpenStudy (anonymous):

of course, where do I write one out

OpenStudy (fibonaccichick666):

Thanks! just click on my name and it should pop up

OpenStudy (anonymous):

Strange I hover over your name and I know the tooltip so prompt me to writing one but its not popping up

OpenStudy (anonymous):

tooltip is supposed to*

OpenStudy (anonymous):

Got it, it stopped acting stupid

OpenStudy (fibonaccichick666):

it takes a minute, it will say something like "hey write a testimonial! they will appreciate it!"

OpenStudy (anonymous):

done :)

OpenStudy (fibonaccichick666):

Thank you and it was wonderful =) Good luck with excel and your classes!

OpenStudy (anonymous):

Thanks a bunch!

OpenStudy (anonymous):

So is it 54 students left after the 12 months or 51?

OpenStudy (fibonaccichick666):

54

OpenStudy (anonymous):

How do type the formula for this problem in excel?